OpenStudy (anonymous):

The population of a city is growing according to the exponential model P = Ce^kt, where P is the population in thousands and t is measures in years. If the population doubles every 11 years what is k, the city's growth rate? Round answer to the nearest hundredth. A) 2.8% B) 4.4% C) 6.3% D) 8.9%

zepdrix (zepdrix):

\[\large P=Ce^{kt}\] At time \(\large t=0\) we'll say that we have a population amount of \(\large P=Ce^0\). Which simplifies to,\[\large P=C\] 11 years later, at time \(\large t=11\), we have double the population that we started with. We ca write it like this,\[\large 2P=Ce^{11k}\]

zepdrix (zepdrix):

So we have a pair of equations. Let's take the second equation and divide it by the first equation.\[\large \frac{2P}{P}=\frac{Ce^{11k}}{C}\]

zepdrix (zepdrix):

From here, you should have some nice cancellations. You'll just need to make sure of the natural log in order to solve for \(\large k\). Think you can solve it from here? :)

OpenStudy (anonymous):

Ah, I think I'll be able to now! Thank you very much! :)